# A Frank-Wolfe-based primal heuristic for quadratic mixed-integer optimization

### v1

@@ 3.4 Convexification of quadratic binary problems
where $Q\in\mathbb{R}^{n\times n}$ is a symmetric matrix and $d\in\mathbb{R}^{n}$ is a linear term. Since all variables are binary, we can leverage the identity $x_{i}^{2}=x_{i}$ to reformulate the problem and “increase” convexity.
Formally, we replace $(Q,d)$ with $(Q_{\ell},d_{\ell})$, such that the convexified matrix $Q_{\ell}$ has a proportion $\ell$ of nonnegative eigenvalues, where $0\leqslant\ell\leqslant 1$. This is achieved through spectral decomposition of $Q$, giving its eigenvalues $\lambda_{1},\dots,\lambda_{n}$ from which we can easily obtain the index $i$ such that $Q_{\ell}=Q+\lambda_{i}I$ satisfies the desired property. Accordingly, $d$ is mapped onto $d_{\ell}=d-\lambda_{i}/2$ to preserve the original objective value at all binary points.
This partial convexification balances problem tractability and the original problem structure, making convexification a tunable process controlled by the parameter $\ell$. A whole stream of work studies hardness [30] and algorithms [26] for quadratic optimization when a small number of eigenvalues are negative. Given the finite time limit, there is indeed a trade-off between the deformation of the initial objective function, whose Lipschitz constant is strongly affected by the transformation, and the advantages of mitigating nonconvexity. In practice, we explore different parameters $\ell$ in parallel on multiple threads. Computational experiments (see Section 5.6) show that some convexification indeed helps but also confirm that full convexification can be detrimental to the search for primal solutions since the original objective is dominated by the convexifying term, sometimes by orders of magnitude, .
@@
The tightest possible convexification can be obtained then via the SDP that finds the optimal vector $u$ such that $Q-2\operatorname{diag}(u)$ is positive semidefinite while minimizing the linear term. This can be formulated as:
@@
We did not explore this approach though as we were primarily interested in finding primal solutions and the tightness of the convexification was not critical to the empirical performance of our approach, so that the additional computational cost of solving the SDP was not justified.
4 Heuristics
Rounding.
@@ 4 Heuristics
One of the simplest techniques is standard rounding, where fractional values are rounded to the nearest integer. Additionally, Boscia employs probability rounding for binary variables, where variables are randomly fixed to 0 or 1 based on probability distributions, and Frank-Wolfe is then used to solve for the remaining continuous variables. Another specialized rounding method is designed for 0/1 polytopes, ensuring that the rounded solution remains within the feasible region. Furthermore, for problems with simplex-like feasible regions, Boscia implements a specialized rounding heuristic that is aware of the defining hyperplane structure, improving feasibility and efficiency in these cases.
Gradient-based heuristics.
@@
Beyond these basic rounding strategies, we utilize gradient-based heuristics, which explore a set of promising vertices, e.g., a follow-the-gradient heuristic inspired by the chasing gradients paradigm [13]. These heuristics are particularly suitable for Boscia even outside of the competition context since the algorithmic components they require (gradient and linear minimization oracles) are precisely those that are available. The method assumes an LMO-compatible feasible region; in the context of the competition, this means in particular that nonlinear constraints have been transformed to penalties. The heuristic consists in starting from any point in the feasible region, computed for instance from the LMO with an arbitrary direction. From that point, the heuristic will compute the gradient, and call the LMO to compute another extreme point. The algorithm iterates for a certain budget of maximum iterations or until it cycles to a vertex already encountered. Follow-the-gradient can be viewed as a (nonconvergent) Frank-Wolfe algorithm with a unit step size. Interestingly, the approximation guarantees of our follow-the-gradient heuristics matches the lower bounds of [1]; see also [23] for a proximity result that provides good intuition why such a heuristic can be often powerful. Preliminary computational experiments highlighted that gradient-based heuristics are not essential to performance of our overall framework; we therefore not report results for this heuristic in the main text. The reason is likely that the solutions it obtains are redundant with the ones found during the execution of Frank-Wolfe. This heuristic could however prove useful when using Boscia as an exact solver instead, with the advantage of only assuming access to the constraints as a linear minimization oracle.
Large Neighborhood Search.
Moreover, we combine our approach with the following large neighborhood search heuristics.
  1. 1.
    We introduce ASENS (Active Set Enforced Neighborhood Search), an adaptation of the RENS heuristic [3] that leverages integer vertices sampled during the execution of Frank-Wolfe algorithms. The active set refers to the set of extreme points of the feasible region whose convex combination forms the current iterate. Note that in recent variants of Frank-Wolfe, the active set size is kept small by performing correction steps and favoring steps that are local to the active set (see [33] for the blended pairwise conditional gradients variant and [7] for explanations and convergence guarantees when favoring local steps). In our case, ASENS is used only when more than 50% of variables have the same values across all points in the active set. ASENS then fixes these variables to these values and solves the subproblem. In most cases where ASENS is employed, the subproblem fixes significantly more than 50% of all variables, and the domains of the continuous variables are also reduced to the convex hull of the active set.
  2. 2.
@@
    We implement the undercover heuristic [5], specifically designed for nonlinear problems. Undercover (UC) identifies and fixes a subset of variables such that the resulting subproblem becomes linear and in many cases easier to solve. Our implementation solves a minimum vertex cover integer linear program to determine this subset, then fixes these variables to values from a reference solution obtained from Boscia’s branch-and-bound tree. We also experimented with undercover started from multiple covers that were different enough from each other but did not observe performance improvements.
  3. 3.
@@
    We also incorporate the solution improvement heuristic RINS (Relaxation Induced Neighborhood Search) [14] improvement heuristic, which identifies promising search regions by fixing all variables that have identical values in both the incumbent solution and the current fractional solution of the branch-and-bound tree. Also, RINS is used only when more than 50% of variables have the same values across the incumbent solution and the current fractional solution.
Specialized heuristics for QUBOs.
@@
For QUBOs whose objective matrix has a bipartite underlying graph, an initial solution can be improved on by alternately optimizing over each component of the graph. This heuristic can seem redundant in comparison with the more general approach described above. However, the implementation of this specific case can be made very fast by exploiting this bipartite structure at a very low level in the code. The inspiration for this fast implementation comes from similar problems encountered in quantum communications, namely, the computation of the local bound of a (bipartite) Bell inequality [15]. In practice, however, this approach was overwhelmingly outperformed by the other heuristics described above.
@@ 5 Computational experiments
In this section, we want to evaluate the performance of our heuristic and the impact of its different components. We want to answer the following questions:
@@
We measure performance with different metrics, including the gap at the end of the time limit (compared to a reference dual bound), primal integral (PI), and whether a solution was found altogether. These metrics highlight different important criteria that can be relevant depending on the context.
@@ 5.1 Setup
Evaluation Metrics
To measure the performance of our heuristic, we employ standard metrics, such as the time to first feasible solution, primal gap, and primal integral [4]. The time to first feasible solution (TTF) is the time taken to find the first feasible solution during the search process. The primal gap (Gap) of a heuristic solution $\tilde{x}$ with respect to the best known solution $x^{*}$ is defined as
$\gamma(\tilde{x})=\begin{cases}0&\text{if }|\tilde{x}|=|x^{*}|=0,\\ 1&\text{if }\tilde{x}\cdot x^{*}<0\text{ or no solution is found},\\ \dfrac{|\tilde{x}-x^{*}|}{\max(|\tilde{x}|,|x^{*}|)}&\text{otherwise}.\end{cases}$
The primal integral (PI) captures the evolution of the primal gap over time. Let $t_{0}=0$ be the start of the search process, $t_{1},\ldots,t_{s-1}$ be the points in time when a new incumbent solution is found, and $t_{s}=T$ be the end of the search process, then the primal integral is defined as
$P(T)=\sum_{i=1}^{s}\gamma(\tilde{x}_{i})\cdot(t_{i}-t_{i-1}).$
@@
Parallelization Strategy
Our parallel implementation strategy executes our algorithm using varying configurations in each thread. Specifically, we ran our heuristics with different power penalty parameters $p$ ranging from 1.2 to 1.8. For problems with binary quadratic objectives, we employed various convexification strategies to explore diverse solution approaches simultaneously. The choice of different values for $p$ and convexification parameters is motivated by preliminary experiments that indicate there is no single best value for these parameters across all instances (see Section 5.6) and trying different values allows us to explore different regions of the solution space.
@@
5.2 Results
In Table 1 we report our finding on QPLIB and the two categories of instances: MIQCQP and MIQP (without quadratic constraints). For each category we report the following metrics: the number of instances for which we found a feasible solution, and for instances where we found a solution, we report the average time to the first feasible solution, the optimality gap compared to the best known solution value from QPLIB, and the primal integral (PI) also using the best known solution as a reference.
@@ 5.2 Results
Category Found TTF Gap PI
MIQP 163/166 1.76 4.06 5.23
MIQCQP 98/153 7.31 25.30 50.56
All 261/319 3.18 11.57 12.77
Our heuristic is particularly effective for MIQP instances, and is able to find feasible solutions for all but three instances in this category and achieves an average optimality gap of 4.06% on instances where we found a solution. Notably, we achieved optimality (0% gap) on 104 instances. The MIQCQP category is more challenging, and we are able to find feasible solutions for 98 out of 153 instances, with an average optimality gap of 25.30%. We achieved optimality on 28 instances. In total we found feasible solutions for 261 out of 319 instances, with an average time to first feasible solution of 3.18 seconds. The average optimality gap was 11.57%, with the average primal integral at 12.77.
@@ 5.4 Impact of LNS heuristics
To assess the impact of the LNS heuristics finding feasible solutions, we compare the performance of the base heuristic with the LNS-off variant, which does not use any LNS heuristics, the ASENS-only variant, which only uses the ASENS heuristic, and the UC-only variant, which only uses the undercover heuristic. The results are summarized in Table 3 and demonstrate both the significant contribution of the LNS heuristics to our framework’s performance and the robustness of the underlying approach without them.
In particular, the LNS-off setting, which relies exclusively on LMO calls without any quadratic subproblem solving, successfully finds feasible solutions for a substantial amount of 226 instances, with an average optimality gap of 28.67% and a primal integral of 29.42 over the 258 instances where at least one of the settings found a solution. This result demonstrates that our framework’s core mechanism of leveraging Frank-Wolfe iterations over the convex hull of integer-feasible points is robust even when restricted to LMO calls. Similarly, the UC-only setting performs well, finding feasible solutions for 249 instances, also without solving any quadratic subproblems. The ASENS-only setting leads to an increased number of feasible solutions found, when compared to LNS-off, and achieves the closest performance with respect to the primal integral and optimality gap to the base heuristic. As shown in Table 3 and more detailed in Figure 3, the contribution of the LNS heuristics is particularly evident in the MIQCQP category, where the base heuristic finds feasible solutions for 92 instances, while the LNS-off setting finds solutions for only 74 instances, and the ASENS-only and UC-only settings find solutions for 86 and 85 instances, respectively.

### v2

@@ 3.4 Convexification of quadratic binary problems
where $Q\in\mathbb{R}^{n\times n}$ is a symmetric matrix and $d\in\mathbb{R}^{n}$ is a linear term. Since all variables are binary, we can use the identity $x_{i}^{2}=x_{i}$ to reformulate the problem and “increase” convexity.
Formally, we replace $(Q,d)$ with $(Q_{\ell},d_{\ell})$, such that the convexified matrix $Q_{\ell}$ has a proportion $\ell$ of nonnegative eigenvalues, where $0\leqslant\ell\leqslant 1$. This is achieved through a spectral decomposition of $Q$, giving its eigenvalues $\lambda_{1},\dots,\lambda_{n}$ from which we can easily obtain the index $i$ such that $Q_{\ell}=Q+\lambda_{i}I$ satisfies the desired property. Accordingly, $d$ is mapped onto $d_{\ell}=d-\lambda_{i}/2$ to preserve the original objective value at all binary points.
This partial convexification balances problem tractability and the original problem structure, making convexification a tunable process controlled by the parameter $\ell$. A whole stream of work studies hardness [34] and algorithms [29] for quadratic optimization when a small number of eigenvalues are negative. Given the finite time limit, there is indeed a trade-off between the deformation of the initial objective function, whose Lipschitz constant is strongly affected by the transformation, and the advantages of mitigating nonconvexity. In practice, we explore different parameters $\ell$ in parallel on multiple threads. Computational experiments (see Section 6.6) show that some convexification indeed helps but also confirm that full convexification can be detrimental to the search for primal solutions since the original objective is dominated by the convexifying term, sometimes by orders of magnitude.
@@
The tightest possible convexification can then be obtained via the SDP that finds the optimal vector $u$ such that $Q-2\operatorname{diag}(u)$ is positive semidefinite while minimizing the objective value. This can be formulated as:
@@
We did not explore this approach, though, as we were primarily interested in finding primal solutions and the tightness of the convexification was not critical to the empirical performance of our approach, so that the additional computational cost of solving the SDP was not justified.
4 Primal Heuristics
Our framework incorporates several primal heuristics to find feasible solutions from points generated during the Frank-Wolfe iterations.
4.1 Rounding
@@
One of the simplest techniques is standard rounding, where fractional values are rounded to the nearest integer. Additionally, Boscia employs probability rounding for binary variables, where variables are randomly fixed to 0 or 1 using the entries as probabilities, and Frank-Wolfe is then used to solve for the remaining continuous variables. We present in Algorithm 3 the pseudo-code for this rounding technique. Another specialized rounding method is designed for 0/1 polytopes, ensuring that the rounded solution remains within the feasible region. Furthermore, for problems with simplex-like feasible regions, Boscia implements a specialized rounding heuristic that is aware of the defining hyperplane structure, improving feasibility and efficiency in these cases. Simplex-like feasible regions are the (scaled) unit simplex and the (scaled) probability simplex.
Algorithm 3 Probability Rounding for a Mixed-Binary Program
1:fractional point $x$, set of integer variables II
2:for $i\in I$ do
3:  $v\;\sim\;\text{Bernoulli}(x_{i})$
4:  if $v=1$ then
5:   $\hat{x}_{i}\leftarrow\min\{1,\lceil x_{i}\rceil\}$
6:  else
7:   $\hat{x}_{i}\leftarrow\max\{0,\lfloor x_{i}\rfloor\}$
8:  end if
9:end for
[…]
4.2 Gradient-based heuristics
@@
Beyond the rounding strategies described above, we utilize gradient-based heuristics, which explore a set of promising vertices, e.g., a follow-the-gradient heuristic inspired by the chasing gradients paradigm [15]. These heuristics are particularly suitable for Boscia since the algorithmic components they require (gradient and linear minimization oracles) are precisely those that are available. The method assumes an LMO-compatible feasible region; in the context of quadratic programming, this means in particular that quadratic constraints have been transformed to objective penalties. The heuristic consists in starting from any point in the feasible region, computed for instance from the LMO with an arbitrary direction. From that point, the heuristic will compute the gradient, and call the LMO to compute another extreme point. The algorithm iterates for a certain budget of maximum iterations or until it cycles to a vertex already encountered. Follow-the-gradient can be viewed as a (nonconvergent) Frank-Wolfe algorithm with a unit step size. Interestingly, the approximation guarantees of our follow-the-gradient heuristics matches the lower bounds of [1]; see also [26] for a proximity result that provides good intuition why such a heuristic can be often powerful. Preliminary computational experiments highlighted that gradient-based heuristics are not essential to the performance of our overall framework, therefore we do not report results for this heuristic in the main text. The reason is likely that the solutions it obtains are redundant with the ones found during the execution of Frank-Wolfe. This heuristic could, however, prove useful when using Boscia as an exact solver instead, with the advantage of only assuming access to the constraints as a linear minimization oracle, or also as seeding for large neighborhood search heuristics (see next section).
4.3 Large neighborhood search
Large Neighborhood Search (LNS) heuristics are powerful techniques for mixed-integer optimization. Many well-known LNS heuristics for MIP rely on LP relaxation solutions [3, 16], or reference solutions generated in other components such as cutting planes [12] or other primal heuristics [31]. In all these approaches, the reference solutions are fractional and feasible for the LP relaxation. We propose instead to leverage the current Frank–Wolfe iterate and the integer-feasible vertices sampled during the execution of the Frank-Wolfe algorithm within the branch-and-bound process. Note that these vertices are not necessarily feasible for the original MIQCQP continuous relaxation since we relax quadratic constraints via penalties.
Active Set Enforced Neighborhood Search.
Algorithm 4 introduces Active Set Enforced Neighborhood Search (ASENS), a new heuristic that capitalizes on the structure of the FW active set by searching in the neighborhood spanned by the active set points. Recent FW variants maintain small active sets through correction and local steps (see [37] for the blended pairwise conditional gradients variant and [7] for guarantees when favoring local steps). We apply ASENS only when more than 50% of the variables take the same value across all active set vertices; ASENS then fixes these variables accordingly and solves the resulting subproblem. In practice, the subproblem often fixes substantially more than half of the variables, and we also restrict the continuous variables to the convex hull of the active set. Preliminary experiments showed that follow-the-gradient is often redundant with iterates encountered during FW itself, while ASENS provides complementary improvements.
Algorithm 4 Active Set Enforced Neighborhood Search (ASENS)
1:active set vertices $S=\{v^{1},\dots,v^{t}\}$, MIQCQP as in (1), threshold $\tau\in(0,1)$
2:Let $J=\{j:\;v^{1}_{j}=\dots=v^{t}_{j}\}$
3:if $|J|/n<\tau$ then $\triangleright$ not enough variables fixed
4:  return
5:end if
6:Fix $x_{j}\leftarrow v^{(1)}_{j}$ for all $j\in J$
7:Restrict bounds of $x_{j}$ to $\min_{i}v^{i}_{j},\max_{i}v^{i}_{j}]$ for all $j\notin J$
8:Solve the resulting MIQCQP
Undercover.
@@ 4.3 Large neighborhood search
We implement a variant of the undercover heuristic [5], specifically designed for nonlinear problems. Undercover (UC) identifies and fixes a subset of variables such that the resulting subproblem becomes linear and, in many cases, easier to solve. Our implementation solves a minimum vertex cover integer linear program to determine this subset, then fixes these variables to values from the current Frank–Wolfe iterate from Boscia’s branch-and-bound tree. We also experimented with undercover initialized from multiple diverse covers but did not observe performance improvements.
RINS.
@@
We implement a Frank–Wolfe-based variant of the Relaxation Induced Neighborhood Search (RINS) [16]. While traditional RINS fixes variables that have identical values in both the incumbent solution and the current relaxation solution, our variant identifies promising search regions by fixing all variables that have identical values in both the incumbent solution and the current Frank–Wolfe iterate from the branch-and-bound tree. Similar to ASENS, RINS is used only when more than 50% of variables agree between the two solutions.
4.4 Specialized heuristics for QUBOs
@@
For QUBOs whose objective matrix has a bipartite underlying graph, an initial solution can be improved on by alternately optimizing over each component of the graph. This heuristic can seem redundant in comparison with the more general approach described above. However, as there is a simple closed-form solution for each step of the alternating optimization, this specific case can be implemented very efficiently. The inspiration for this comes from similar problems encountered in quantum communications, namely, the computation of the local bound of a (bipartite) Bell inequality [17]. In practice, however, this approach was consistently outperformed by the other heuristics described above.
@@ 5 Computational experiments
In this section, we evaluate the performance of our heuristic and the impact of its different components. We seek to answer the following questions:
@@
We measure performance with respect to different metrics, including the time to the first feasible solution (TTF), the gap at the end of the time limit (compared to a reference dual bound), and the primal integral (PI). These metrics highlight different important criteria that can be relevant depending on the context.
@@ 5.1 Setup
Evaluation metrics.
[…]
$\gamma(\tilde{x})=\begin{cases}0&\text{if }|\tilde{f}|=|f^{*}|=0,\\ 1&\text{if }\tilde{f}\cdot f^{*}<0,\,\text{or no solution is found},\\ \dfrac{|\tilde{f}-f^{*}|}{\max(|\tilde{f}|,|f^{*}|)}&\text{otherwise}.\end{cases}$
The primal integral captures the evolution of the primal gap over time. Let $t_{0}=0$ be the start of the search process, $t_{1},\ldots,t_{s-1}$ be the points in time when a new incumbent solution is found, and $t_{s}=T$ be the end of the search process, then the primal integral is defined as
$P(T)=\sum_{i=1}^{s}\gamma(\tilde{x}^{i})\cdot(t_{i}-t_{i-1}).$
@@
Parallelization strategy.
Our parallel implementation strategy executes our algorithm using varying configurations in each thread. Specifically, we ran our heuristics with different power penalty parameters $p$ ranging from 1.2 to 1.8. For problems with binary quadratic objectives, we employed various convexification strategies to explore diverse solution approaches simultaneously. The choice of different values for $p$ and convexification parameters is motivated by preliminary experiments that indicate there is no single best value for these parameters across all instances (see Section 6.6), and trying different values allows us to explore different regions of the solution space.
@@
6 Results
In Table 1 we report our finding on QPLIB and the two categories of instances: MIQCQP and MIQP (without quadratic constraints). For each category we report the following metrics: the number of instances for which we found a feasible solution, and for instances where we found a solution, we report the average optimality gap compared to the best known solution value from QPLIB, the primal integral also using the best known solution as a reference, and time to the first feasible solution.
@@ 6 Results
Category Found Gap (%) PI TTF
MIQP 163/166 4.06 5.23 1.76
MIQCQP 98/153 25.30 50.56 7.31
Total 261/319 11.57 12.77 3.18
Our heuristic is particularly effective for MIQP instances, and is able to find feasible solutions for all but three instances in this category, and achieves an average optimality gap of 4.06% on instances where we found a solution. Notably, we achieved optimality (0% gap) on 104 instances. The MIQCQP category is more challenging, and we are able to find feasible solutions for 98 out of 153 instances, with an average optimality gap of 25.30%. We achieved optimality on 28 instances. In total we found feasible solutions for 261 out of 319 instances, with an average time to first feasible solution of 3.18 seconds. The average optimality gap was 11.57%, with the average primal integral at 12.77.
6.1 Competition and improvements on QPLIB
Table 2: Improved solutions for QPLIB instances. The columns indicate the instance name, the objective sense, the previous best-known solution value, the new solution value found by our heuristic, and the relative improvement (gap) in percentage.
Instance Obj. Sense Previous Best New Solution Gap (%)
QPLIB_2169 max 29.0 30.0 3.45
QPLIB_2174 max 150.0 152.0 1.33
QPLIB_2205 max 88.0 90.0 2.27
QPLIB_3347 min 3,819,920 3,818,879 0.03
QPLIB_3584 min -25,254 -25,386 0.52
QPLIB_3709 min 5,726,530 5,710,645 0.28
QPLIB_3860 min -19,685 -20,161 2.42
QPLIB_10022 min 6,267,782 1,374,066 78.07
6.2 Comparison to state-of-the-art solvers
To assess the quality of our heuristic relative to state-of-the-art solvers, we compare its performance against Gurobi 12.0.1 and the open-source solver SCIP (pre-release version 10.0.0, GitHub commit 30844abadc [^2] https://github.com/scipopt/scip[10] on the QPLIB benchmark instances. Both solvers were run with default settings and a time limit of 300 seconds on the same hardware as our heuristic. Table 3 summarizes the results, comparing the solution quality achieved by each solver within the time limit.
Table 3: Performance comparison of the parallel heuristic against Gurobi and SCIP on QPLIB instances with a 300-second time limit. The table shows the number of instances where the heuristic found a solution with the same, better, or worse objective value compared to each solver. The “None” column indicates instances where no solution was found by either method.
vs. Gurobi vs. SCIP
Type Same Better Worse None Same Better Worse None
MIQP 93 30 43 0 40 94 31 1
MIQCQP 27 1 101 24 15 26 81 31
Our heuristic demonstrates competitive performance, particularly on MIQP instances, where it finds better solutions than Gurobi on 30 instances and better solutions than SCIP on 94 instances. For MIQCQPs, both Gurobi and SCIP outperform our heuristic, highlighting the substantial challenges posed by quadratic constraints.
@@ 6.4 Impact of LNS heuristics
To assess the impact of the LNS heuristics finding feasible solutions, we compare the performance of the base heuristic with the LNS-off variant, which does not use any LNS heuristics, the ASENS-only variant, which only uses the ASENS heuristic, and the UC-only variant, which only uses the undercover heuristic. All variants still use the rounding heuristics described in Section 4.1. The results are summarized in Table 5 and demonstrate both the significant contribution of the LNS heuristics to our framework’s performance and the robustness of the underlying approach without them.
In particular, the LNS-off setting, which relies exclusively on LMO calls without any quadratic subproblem solving, successfully finds feasible solutions for a substantial amount of 226 instances, with an average optimality gap of 28.67% and a primal integral of 29.42 over the 258 instances where at least one of the settings found a solution. This result demonstrates that our framework’s core mechanism of leveraging Frank-Wolfe iterations over the convex hull of integer-feasible points is robust even when restricted to LMO calls. Similarly, the UC-only setting performs well, finding feasible solutions for 249 instances, also without solving any quadratic subproblems. The ASENS-only setting leads to an increased number of feasible solutions found, when compared to LNS-off, and achieves the closest performance with respect to the primal integral and optimality gap to the base heuristic. As shown in Table 5 and in Figure 3, the contribution of the LNS heuristics is particularly evident in the MIQCQP category, where the base heuristic finds feasible solutions for 92 instances, while the LNS-off setting finds solutions for only 74 instances, and the ASENS-only and UC-only settings find solutions for 86 and 85 instances, respectively.